If the total cost is R750 and the average fixed cost is R15 when 30 units of output are
produced, then the average variable cost at that level of output is …

1. R15.
2. R10.
3. R450.
4. Impossible to determine.
1
Expert's answer
2019-02-24T14:38:08-0500

Fixed cost=15*30=R450

Total variable cost=750-450=R300

Average variable cost=300/30=R10

2.R10